OpenStudy (anonymous):

A house painting company charges $376 plus $12 per hour. Another company $280 plus $15 per hour. A. how long is a job for which both companies will charge the same amount? B.What will the cost be?

OpenStudy (tkhunny):

Can you build a linear function for each? House Painting Company h(t) = ?? Other Company O(t) = ??

OpenStudy (texaschic101):

376 + 12h = 280 + 15h 376 - 280 = 15h - 12h 96 = 3h 96/3 = h 32 = h 376 + 12(32) =376 + 384 = 760 280 + 15(32) = 280 + 480 = 760 ANSWER : (a) 32 hours (b) $ 760

OpenStudy (tkhunny):

Gaaa!! I guess we'll never know if samkurk can do it.

OpenStudy (texaschic101):

My bad....I am sorry....I shouldn't have given the answer :(